# A HHO formulation for variable density incompressible flows where the density is purely advected

### v1

@@ 4.3 Rayleigh-Taylor instability
In order apply the ESDIRK-HHO formulation to a realistic variable density incompressible flow problem we consider a low Atwood number Rayleigh-Taylor Instability (RTI) configuration. The test case, originally proposed by Tryggvason [24], is setup as proposed by Guermond et al. [16].
@@
We consider the time evolution of two layers of immiscible fluids initially at rest in the rectangular domain $\Omega=(-\frac{d}{2},\frac{d}{2})\times(-2d,2d)$, where $d$ is the reference length. Evolution of the flow field is driven by the downward gravity field as the heavy and the light fluids occupy the top and the bottom layers, respectively. Denoting by $\rho_{\mathrm{t}}$ and $\rho_{\mathrm{b}}$ the density of the top and bottom fluids, respectively, the density ratio $\frac{\rho_{\rm{t}}}{\rho_{\rm{b}}}$ is set to 3, hence the Atwood number, defined as $\mathrm{At}=(\rho_{\mathrm{t}}-\rho_{\mathrm{b}})/(\rho_{\mathrm{t}}+\rho_{\mathrm{b}})$, is equal to one half [24]. Following [16], a smooth transition between the two fluids is guaranteed relying on the formula
$\frac{\rho(\bm{x},t=0)}{\rho_{\mathrm{b}}}=2+\mathrm{tanh}\left(\frac{\bm{x}\cdot\bm{j}-\eta(\bm{x})}{0.01d}\right),$ (22)
@@
where $\eta(\bm{x})=-0.1d\,\mathrm{cos}\left(\frac{2\pi}{d}\bm{x}\cdot\bm{i}\right)$ is the initial position of the perturbed interface between the two fluids. The unit vectors $\bm{i}$ and $\bm{j}$ have horizontal orientation pointing right and vertical orientation pointing up, respectively. The $L^{2}$ projection of the analytical density distribution in Eq. (22) provides the initial density field and the source term in the momentum equation reads $\bm{f}=-\rho g\bm{j}$.
@@
The Reynolds number is defined as $\mathrm{Re}=\frac{\rho_{\mathrm{b}}d^{\frac{3}{2}}g^{\frac{1}{2}}}{\mu}$, note that $d^{\frac{1}{2}}g^{\frac{1}{2}}$ is the reference velocity. Assuming the symmetry of the initial condition is maintained during the time evolution, the computational domain has been restricted to $\Omega=(0,\frac{d}{2})\times(-2d,2d)$ and discretized with a triangular elements mesh with $\operatorname{card}(\mathcal{T}_{h})=65~536$. The computational mesh is obtained splitting in two, along the diagonal, the quadrilateral elements of a $64{\times}512$ cartesian grid. Symmetry boundary conditions, implemented as homogeneous Dirichlet and Neumann BCs in the normal and tangential directions, respectively, are imposed on the computational domain boundary $\partial\Omega$. The polynomial degree of the HHO discretization is $k=6$ and time integration is conducted by means of the second order accurate three stages TR-BDF2 ESDIRK scheme, see [23], within the interval $(0,4]$ discretized using 16001~600 uniform time steps.
@@
In practice we consider unit reference lenght, unit gravity and unit $\rho_{\mathrm{b}}$, so that also the reference velocity is equal to one and $\rho_{\mathrm{t}}=3$. Thus, we set $\mu=\frac{1}{\mathrm{Re}}$ and we run computations considering two Reynolds numbers: $\mathrm{Re}{=}1~000$ and $\mathrm{Re}{=}5~000$. The time evolution of the density field is shown in Figures 1 and 2 at times $t_{\mathrm{Trg}}\simeq 1,1.5,1.75,2,2.25,2.5,2.75$ in the time scale of Tryggvason, where $t_{\mathrm{Trg}}=t\sqrt{\mathrm{At}}$. In practice, for a given a value of $t_{\mathrm{Trg}}$, we select the closest available time point among the 16001~600 available by rounding off to the closest integer the quantity $\frac{t_{\mathrm{Trg}}}{\sqrt{\mathrm{At}}\,\delta t}$, where $\delta t=\frac{4}{1~600}$.
The density field at $\mathrm{Re}{=}1~000$ is in very good agreement with the results of Guermond et al. [17]. As opposite, moving to $\mathrm{Re}{=}5~000$, our computations show more wiggles at the interface between the two fluids, especially at early stages of the flow field evolution. Moreover, the peripheral curling filaments interact with the bulk of the descending plume, a phenomena which does not take place in our results of reference [17]. We remark that the HHO-ESDIRK formulation presented in this work lacks spurious oscillations control while, on the contrary, in [17], an entropy viscosity stabilization was employed. Accordingly, it is somehow expected that the highest Reynolds number computation, associated with a sharper interface featuring strong variations of all the relevant physical quantities, is in worse agreement.
Interestingly, the results at $\mathrm{Re}{=}5~000$ are in better agreement with those those obtained with a high-order Discontinuous Galerkin formulation lacking oscillation control in Bassi et al. [3]. The agreement is quite strong for $t\leq 2$ while the latter stages of time evolution show a different behavior. This confirm the statement that an accurate and detailed prediction of the RTI flow for $t\geq 1.5$ and and $\mathrm{Re}\geq 5000$ is a difficult task [16].
@@
Figure 1: RTI at $\mathrm{At}=0.5$ and $\mathrm{Re}=1~000$ . From left to right, evolution of the density field at selected time points $t_{\mathrm{Trg}}=t\sqrt{\mathrm{At}}\simeq 1,1.5,1.75,2,2.25,2.5,2.75$.
@@
Figure 2: RTI at $\mathrm{At}=0.5$ and $\mathrm{Re}=5~000$. From left to right, evolution of the density field at selected time points $t_{\mathrm{Trg}}=t\sqrt{\mathrm{At}}\simeq 1,1.5,1.75,2,2.25,2.5,2.75$.
@@ 5 Conclusions
We devised and numerically validated an ESDIRK-HHO formulation of the incompressible Navier–Stokes equations with variable density that is capable of simulating mixtures of immiscible fluids reaching high-orders of accuracy in both space and time. Interestingly, the role of the pressure at the continuous level is closely replicated at the discrete level. Indeed, in the one hand pressure is the Lagrange multiplier of the incompressibilty constraint and, on the other hand, the hybrid pressure space enforces $\bm{H}(\operatorname{div})$-conformity starting from non-conforming polynomial spaces. The distinguishing features, that is pressure-robustness, exact incompressibility and pure advection of the density, have been numerically demonstrated.
Since the spatial discretization requires primitive variables in order to achieve $\bm{H}(\operatorname{div})$-conformity, while the time derivatives are formulated in terms of conservative variables, the ESDIRK time marching strategy has been carefully designed. The resulting scheme is fully implicit and fully hybrid, and static condensation of element unknowns has been employed to alleviate the memory footprint of the Jacobian matrix. To further improve the performance of iterative linear solvers a $p$-multigrid preconditioner has been fruitfully exploited when $k>0$. Note that, thanks to the use of non-conforming polynomial spaces, inherited coarse grid operators are trivial to construct. The ability of simulating realistic flow configurations is demonstrated tackling the Rayleigh-Taylor instability. Notably, high-order computations are carried out without oscillation control and the evolution of the density field is in good agreement with previously published works.
Further research efforts are required to target physically relevant multi-phase flows and higher Atwood numbers flow configurations. In particular, future works will consider the extension of the HHO-ESDIRK formulation to the Cahn-Hilliard-Navier-Stokes (CHNS) system.

### v3

@@ 4.3 Rayleigh-Taylor instability
In order apply the ESDIRK-HHO formulation to realistic variable density incompressible flow problems we consider a low and a high Atwood number Rayleigh-Taylor Instability (RTI). The test case, originally proposed by Tryggvason [26], is set up as proposed by Guermond et al. [17].
@@
We consider the time evolution of two layers of immiscible fluids initially at rest in the rectangular domain $\Omega=(-\frac{d}{2},\frac{d}{2})\times(-2d,2d)$, where $d$ is the reference length. Evolution of the flow field is driven by the downward gravity field as the heavy fluid lays on top of the light fluid. The density ratio $\frac{\rho_{\rm{t}}}{\rho_{\rm{b}}}$ is set to three and seven, where $\rho_{\mathrm{t}}$ and $\rho_{\mathrm{b}}{=}1$ are the density of the top and bottom fluids, respectively. Accordingly, the Atwood number, defined as $\mathrm{At}=(\rho_{\mathrm{t}}-\rho_{\mathrm{b}})/(\rho_{\mathrm{t}}+\rho_{\mathrm{b}})$, is ${\rm lAt}{=}0.5$ (low Atwood) and ${\rm hAt}{=}0.75$ (high Atwood). Following [17], a smooth transition between the two fluids is guaranteed relying on the following expressions for the initialization of the density field
[…]
@@
where $\eta$ and $\gamma$ define the initial position of the perturbed interface in the low and the high Atwood configuration, respectively. The unit vectors $\bm{i}$ and $\bm{j}$ have horizontal orientation pointing right and vertical orientation pointing up, respectively. The $L^{2}$ projection of the analytical density distribution in Eq. (33) provides the initial density field and the source term in the momentum equation reads $\bm{f}=-\rho g\bm{j}$.
@@
Assuming the symmetry of the initial condition is maintained during the time evolution, the computational domain has been restricted to $\Omega=(0,\frac{d}{2})\times(-2d,2d)$. We consider two uniform triangular elements grids obtained splitting in two, along the diagonal, the quadrilateral elements of a $32{\times}256$ and a $256{\times}2048$ Cartesian grid, resulting in $\operatorname{card}(\mathcal{T}_{h})=16~384$ and $\operatorname{card}(\mathcal{T}_{h})=1~048~576$ for the coarse and the fine grid, respectively. Symmetry boundary conditions, implemented as homogeneous Dirichlet in the normal direction and homogeneous Neumann in the tangential direction, are imposed on the computational domain boundary $\partial\Omega$. Time integration is conducted by means of the second order accurate three stages TR-BDF2 ESDIRK scheme, see [25], using 32003~200 uniform time steps. The time integration intervals are $(0,4]$ and $(0,3.75)$, respectively, in the low and the high Atwood number configuration.
4.3.1 Low Atwood number computations
In this section we present and discuss the numerical results in the low Atwood number configuration focusing on the evolution of the density field. The computations are performed with both the HHO-$\pi^{k}$ and the HHO-$\pi^{k+1}$ formulations, the polynomial degree is $k=6$ for the coarse grid and $k=1$ for the fine grid. We remark that, even if $k>0$ discretizations are not bounds preserving due to the onset of spurious oscillations at sharp interfaces, the moderate density ratio associated with the low Atwood number setup allows to disregard oscillation control strategies.
@@ 4.3.1 Low Atwood number computations
The Reynolds number is defined as $\mathrm{Re}=\frac{\rho_{\mathrm{b}}d^{\frac{3}{2}}g^{\frac{1}{2}}}{\mu}$, note that $d^{\frac{1}{2}}g^{\frac{1}{2}}$ is the reference velocity. In practice we consider unit reference length, unit gravity and unit $\rho_{\mathrm{b}}$, so that also the reference velocity is equal to one. Thus, we set $\mu=\frac{1}{\mathrm{Re}}$ and we run computations considering two Reynolds numbers: $\mathrm{Re}{=}1~000$ and $\mathrm{Re}{=}5~000$. The time evolution of the density field for HHO-$\pi^{k}$ is shown in Figures 1 and 2 at times $t_{\mathrm{Trg}}\simeq 1,1.5,1.75,2,2.25,2.5,2.75$ in the time scale of Tryggvason, where $t_{\mathrm{Trg}}=t\sqrt{\mathrm{At}}$. In practice, for a given a value of $t_{\mathrm{Trg}}$, we select the closest available time point among the 32003~200 available by rounding off to the closest integer the quantity $\frac{t_{\mathrm{Trg}}}{\sqrt{\mathrm{At}}\,\delta t}$, where $\delta t=\frac{t_{F}}{3~200}$.
At $\mathrm{Re}{=}1~000$ the density fields provided by k=1k{=}1 HHO-$\pi^{k}$ over the fine grid and k=6k{=}6 HHO-$\pi^{k}$ over the coarse grid are closely matched and the results are in very good agreement with those obtained by Guermond et al. [18]. As opposite, at $\mathrm{Re}{=}5~000$, some differences can be appreciated between the density fields provided by k=1k{=}1 and k=6k{=}6 approximations. In particular, k=6k{=}6 results show early breakup of the spiraling filament while k=1k{=}1 results show interface sprouts in the proximity of the computational domain boundaries, especially at later stages of the flow field evolution. Moreover, at k=6k{=}6, the peripheral curling filaments interact with the bulk of the descending plume, a phenomena which does not take place in our results of reference [18], where an entropy viscosity stabilization was employed. Since high Reynolds number computations are associated with sharper interfaces featuring strong variations of all relevant physical quantities, the aforementioned discrepancies can be ascribed to the lack of oscillation control. In order to highlight the benefits of $p$-multigrid solutions strategies combined with static condensation, it is worth to mention that k=6k{=}6 HHO over the coarse grid provides a fourteen fold decrease of the global Jacobian matrix dimension compared to k=1k{=}1 HHO over the fine grid, leading to significant advantages in terms of computation time.
A comparison between HHO-$\pi^{k}$ and HHO-$\pi^{k+1}$ is reported in Figure 3 and Figure 4, focusing on the position of the interface between the two fluids. Accordingly, by setting the threshold to the average density ($\rho{=}2$), we are able to distinguish between the heavy fluid, depicted in white, and the light fluid, depicted in black. To highlight the differences between high-order and low-order accurate results, we depict the whole computational domain $\Omega=(-\frac{d}{2},\frac{d}{2})\times(-2d,2d)$ by gluing together k=6k{=}6 computations rotated by 180 degrees, as if $\Omega^{k=6}=(-\frac{d}{2},0)\times(-2d,2d)$ was the computational domain, with k=1k{=}1 computations over $\Omega^{k=1}=(0,\frac{d}{2})\times(-2d,2d)$, i.e., $\Omega=\Omega^{k=6}\bigcup\Omega^{k=1}$.
At both Re=1000Re{=}1~000 and Re=5000Re{=}5~000, HHO-$\pi^{k}$ and HHO-$\pi^{k+1}$ computations are in very good agreement when $k=1$, while some differences can be appreciated comparing $k=6$ results at Re=5000Re{=}5~000. As previously mentioned, we can conclude that spurious oscillations influence the most the flow field evolution when the interface is sharp and high-degree polynomials are considered. This confirms the statement that an accurate and detailed prediction of the RTI flow for $t\geq 1.5$ and and $\mathrm{Re}\geq 5000$ is a difficult task, as observed by Guermond et al. [17].
@@
Figure 1: RTI at $\mathrm{At}{=}0.5$ and $\mathrm{Re}{=}1~000$ with HHO-$\pi^{k}$. From left to right, evolution of the density field at selected time points $t_{\mathrm{Trg}}=t\sqrt{\mathrm{At}}\simeq 1,1.5,1.75,2,2.25,2.5,2.75$. Top and bottom row, results for k=1k{=}1 over fine grid and k=6k{=}6 over the coarse grid, respectively.
Figure 2: RTI at $\mathrm{At}{=}0.5$ and $\mathrm{Re}{=}5~000$ with HHO-$\pi^{k}$. From left to right, evolution of the density field at selected time points $t_{\mathrm{Trg}}=t\sqrt{\mathrm{At}}\simeq 1,1.5,1.75,2,2.25,2.5,2.75$. Top and bottom row, results for $k=1$ over the fine grid and $k=6$ over the coarse grid, respectively.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: RTI plume close-up view at $\mathrm{At}{=}0.5$ and $\mathrm{Re}{=}1~000$ with HHO-$\pi^{k}$ (top row) and HHO-$\pi^{k+1}$ (bottom row). From left to right, evolution of the interface at selected time points $t_{\mathrm{Trg}}=t\sqrt{\mathrm{At}}\simeq 1.75,2,2.25,2.5,2.75$. In each frame, results obtained with k=6k{=}6 over the coarse grid and k=1k{=}1 over the fine grid results stands on the left and the right, respectively, with respect to a vertical center-line.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: RTI plume close-up view at $\mathrm{At}{=}0.5$ and $\mathrm{Re}{=}5~000$ with HHO-$\pi^{k}$ (top row) and HHO-$\pi^{k+1}$ (bottom row). From left to right, evolution of the interface at selected time points $t_{\mathrm{Trg}}=t\sqrt{\mathrm{At}}\simeq 1.75,2,2.25,2.5,2.75$. In each frame, results obtained with k=6k{=}6 over the coarse grid and k=1k{=}1 over the fine grid results stands on the left and the right, respectively, with respect to a vertical center-line.
4.3.2 High Atwood number results
The high density ratio configuration is challenging from the stability viewpoint, indeed, when high-order accurate formulations are employed, spurious oscillation might lead to negative densities and, eventually, singular mass matrices. Accordingly, we consider k=0k{=}0 HHO-$\pi^{k}$ and HHO-$\pi^{k+1}$ discretizations over the fine one million elements grid. The Reynolds number, defined as $\mathrm{Re}=\frac{\rho_{\mathrm{b}}d^{\frac{3}{2}}g^{\frac{1}{2}}}{\mu}$, is $\mathrm{Re}{=}1~000$ and we set $\mu=\frac{1}{\mathrm{Re}}$. The time evolution of the density field is shown in Figure 5 at times $t\simeq 1,1.5,2,2.5,3,3.5,3.75$. In practice, for a given a value of $t$, we select the closest available time point among the 32003~200 available by rounding off to the closest integer the quantity $\frac{t}{\delta t}$, where $\delta t=\frac{t_{F}}{3~200}$.
We remark that both HHO-$\pi^{k}$ and HHO-$\pi^{k+1}$ are bounds preserving at k=0k{=}0, meaning that the maximum and the minimum density values observed all along the density field evolution correspond to the density of the heavy ($\rho=7$) and the light ($\rho=1$) fluid. Comparing with the reference results of [18], it is clear that the spatial accuracy is not sufficient to properly capture the evolution of the sharp interface between the two fluids. We remark that our computations are in better agreement with the earlier results of [17] where a coarser computational mesh was employed as compared with [18]. Even if the tiniest flow features are missing or misplaced, the position and shape of the main plume is properly replicated. Note that HHO-$\pi^{k+1}$ seems to introduce less numerical dissipation as compared with HHO-$\pi^{k}$, confirming the better convergence rates observed in Section 4.1.
Figure 5: RTI at $\mathrm{At}{=}0.75$ and $\mathrm{Re}{=}1~000$. Top and bottom row, k=0k{=}0 HHO-$\pi^{k}$ and HHO-$\pi^{k+1}$ computations, respectively. From left to right, evolution of the density field at selected time points $t\simeq 1,1.5,2,2.5,3,3.5,3.75$.
@@ 5 Conclusions
We devised and numerically validated an ESDIRK-HHO formulation of the incompressible Navier–Stokes equations with variable density that is capable of simulating mixtures of immiscible fluids reaching high-orders of accuracy in both space and time. Interestingly, the role of the pressure at the continuous level is closely replicated at the discrete level. Indeed, in the one hand pressure is the Lagrange multiplier of the incompressibilty constraint and, on the other hand, the hybrid pressure space enforces $\bm{H}(\operatorname{div})$-conformity starting from non-conforming polynomial spaces. The variant of the method named HHO-$\pi^{k}$ is provably stable thanks to the use of skew-symmetric forms for the transport of mass and momentum. The distinguishing features, that is pressure-robustness, exact incompressibility and pure advection of the density, have been numerically demonstrated.
Since the spatial discretization requires primitive variables in order to achieve $\bm{H}(\operatorname{div})$-conformity, while the time derivatives are formulated in terms of conservative variables, the ESDIRK time marching strategy has been carefully designed. The resulting scheme is fully implicit and fully hybrid, and static condensation of element unknowns has been employed to alleviate the memory footprint of the Jacobian matrix. To further improve the performance of iterative linear solvers a $p$-multigrid preconditioner has been fruitfully exploited when $k>0$. Note that, thanks to the use of non-conforming polynomial spaces, inherited coarse grid operators are trivial to construct.
The ability of simulating realistic flow configurations is demonstrated tackling the Rayleigh-Taylor instability at different Atwood and Reynolds numbers. In the low-Atwood number number setup, the evolution of the density field is obtained with both high-order accurate ($k=6$) formulations on coarse meshes and low-polynomial degree ($k=1$) expansions on fine meshes. Notably, for $\mathrm{At}=0.5$ and $\mathrm{Re}=1000$, the numerical solutions are in very good agreement despite the major difference in degrees-of-freedom count, resulting in significant advantages in terms of computational expense for high-order accurate computations. High-Atwood number computations demonstrate the HHO-ESDIRK formulation is bounds preserving when considering the lowest polynomial degree ($k=0$), which implies that high density ratios can be tackled without compromising stability.
Further research efforts are required to target physically relevant multi-phase flows, in particular, future works will consider the extension of the HHO-ESDIRK formulation to the Cahn-Hilliard-Navier-Stokes (CHNS) system.